OpenStudy (anonymous):

HELP PLEASE One solution is 20% acid and another is 12% acid. How many cubic centimeters of each solution should be mixed to obtain 100 cubic centimeters of a solution that is 16% acid? Show all steps. @timo86m @amistre64

OpenStudy (amistre64):

define 2 equations ... once for volume and the other for percentages

OpenStudy (anonymous):

so you mean..x+y=100 20x +12y=16

OpenStudy (amistre64):

x + y = 100 ,20x + .12y = .16(100)

OpenStudy (amistre64):

close

OpenStudy (amistre64):

1600 is better

OpenStudy (amistre64):

if need be: let x=50+n let y=50 -n

OpenStudy (amistre64):

lol, n=0

OpenStudy (anonymous):

so i should make my second equation 20x + 12y= 1600

OpenStudy (amistre64):

yes

OpenStudy (anonymous):

oh ok.let me work it out and tell u my answer please

OpenStudy (amistre64):

you want .16 of 100; and if you are going to pull out the decimals then thats 16(100)

OpenStudy (anonymous):

ok.let me work it out and tell u what i got

OpenStudy (anonymous):

i had both equations to be 50..@amistre64 , if its right how do i conclude ??

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

x=50 and y=50 thats is

OpenStudy (amistre64):

*thats it

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

thanks very much .